# Demystifying Embedding Spaces using Large Language Models

### v1

@@ 1 Introduction
The success of deep learning has brought forth a paradigm shift in knowledge representation through the concept of embeddings—dense vector representations that capture high-dimensional information about entities, concepts, or relationships in a compact and useful format. Embeddings are ubiquitous, finding application in natural language processing (Mikolov et al., 2013; Devlin et al., 2018), recommender systems (Rendle, 2010), protein sequence modeling (Rives et al., 2021), and more. These embeddings are invaluable, capturing nuanced relationships and semantic structure in data which traditional machine learning (ML) approaches often miss.
Nevertheless, understanding these abstract representations remains challenging. By design, the structure and underlying information carried by an embedding is heavily adapted to the idiosyncrasies of the downstream task, posing a substantial challenge to its interpretation and manipulation. Previous work on ML interpretability offers various task-independent means to interpret embeddings, including dimensionality reduction techniques (e.g., t-SNE (Van der Maaten & Hinton, 2008), UMAP (McInnes et al., 2018)) or concept activation vectors (CAVs) (Kim et al., 2018). While very useful, these techniques are fairly narrow in scope.
As an alternative to such interpretability methods, suppose one could engage with embeddings using natural language to query information not directly expressed by the name or description of the underlying entity/concept. Perhaps one could even extract information from the embedding representations of non-existent or hypothetical entities. Imagine, for instance, an embedding space representing items from an online commerce site, trained using a large corpus of user ratings or purchases, reviews, and other multi-faceted data sources. The embedding representation of an item may implicitly embody intricate details about its quality, usability, design, customer satisfaction, etc. Moreover, suppose one wanted to understand the properties of a hypothetical item at a specific point in the embedding space, say, if we predicted a potential market for such a item. Since no such item (or description) exists, asking a conventional language model to describe this embedding point would be futile. However, a large language model (LLM) trained to interpret the embedding representation itself could handle such a task (see Fig. 1 for an illustration).
@@ 2 Integrating Domain Embeddings with Large Language Models
In this section, we outline our problem formulation. We begin with basic background and motivation for our problem, establishing the terminology used throughout the paper.
2.1 Background
We assume a domain embedding which maps entities in some vocabulary (e.g., a set of known users and items) into a latent space , equipped with a distance metric . The embedding vector of any is typically a representation used in some downstream task, and is trained to encompass the latent features of necessary for this task. For example, in a recommender system, collaborative filtering (CF) (Salakhutdinov & Mnih, 2007; Koren et al., 2011) is often used to generate embedding vectors of users and items s.t. the dot product (or cosine similarity) of and is predictive of user ’s affinity for item . In image classification, might generate embedding representations useful for object detection (Zou et al., 2023), while in search/information retrieval, might embed both queries and documents in to measure a document’s relevance to a query (Zuccon et al., 2015; Luan et al., 2021). Our key requirement is that the vocabulary of the domain embedding can be mapped to a linguistically coherent phrase or token sequence (e.g. the title of a particular movie) that is in distribution for a pretrained LLM like PaLM 2 (Google et al., 2023).
We assume a pretrained LLM, , which maps (length ) sequences of language tokens into another. Modern LLMs are composed of two parts, . The first part is an embedding layer, (not to be confused with the domain embeddings ), which maps tokens to their respective token embedding representations (as distinct from the domain embedding). The second part is the the dense model, , which maps a sequence of token embeddings to a sequence of tokens 222In practice, the dense model outputs logits over tokens, which are sampled by a decoding procedure. (e.g., transformer, Vaswani et al. (2017)). We refer to as a “text-only” LLM.
2.2 Embedding Language Model (ELM)
@@
Training Procedure. Given a downstream loss function , we can differentially optimize the model with parameters over the tasks in by solving . Aside from maximum likelihood-based training, we discuss optimizing using reinforcement learning from AI feedback (Lee et al., 2023) in Appendix G.
Training continuous prompts poses challenges due to constraints over the pretrained token embeddings, which may result in convergence to local minima (Liu et al., 2021). Intuitively, as the pretrained embedding layer readily maps to language space, the newly initialized embedding layer may require numerous updates to map to a similar space. To overcome this, we divide training into two stages. In the first stage, we train the adapter on tasks in by keeping all other parameters (, ) frozen. Since is pretrained, the learned first-stage mapping from to improves convergence in the next stage. In the second stage, we fine-tune the full model by training all parameters (, , ). To increase efficiency, we can update a smaller number of parameters in alternating fashion (Hu et al., 2021). We found that two-stage training is essential for convergence of (see Appendix C for further discussion of the two-stage training procedure).
@@ 3.1 Data
Training Data and Tasks. We test two versions of ELM, one that interprets semantic embeddings of movies, and a second which interprets behavioral embeddings of users. For the first model, we use a variety of tasks to test the model’s ability to extract reasonable semantic information from the domain embeddings. More specifically, we construct 24 movie-focused tasks using a pretrained PaLM 2-L model. For each task, we generate training data by prompting the PaLM 2-L with a movie’s title and additional task-specific information (e.g., writing review, listing characteristics, comparing movies, see details below). We then use PaLM 2-L’s generated output as training targets for ELM. As for training inputs, we provide ELM with the same prompts used in PaLM 2-L, where the title of the movie is replaced by its semantic embedding .666The data-generating prompt also instructs PaLM 2-L not to use the movie title in its output. The training prompt removes this instruction. We emphasize that ELM does not receive any information about the movie (including its title), apart from its semantic embedding vector.
Our 24 movie tasks include single movie semantic tasks, such as describing a movie plot or summarizing a movie; single movie subjective tasks, such as writing positive or negative reviews for a movie; and movie pair subjective tasks, such as comparing characteristics of movies. Appendix D provides a complete description of all 24 tasks, including the prompts used, and sample outputs.
Finally, we train an additional model to interpret user behavioral embeddings, by generating user preference profiles. For this task, we generate a textual summary of a user’s preferences as follows: (a) we sample five positively rated and five negatively rated movies for each user in MovieLens 25M; (b) we then prompt PaLM 2-L to describe (in ten bullet points) characteristics of a person who likes the first five movies, but dislikes the second five. This resulting summary is used as training output for ELM on the user-profile generation task. The resulting task is to generate such a user profile from a user’s behavioral embedding (i.e., interpret the domain embedding). Similar to the movie tasks, no other information about the user, apart from its embedding, is provided to ELM. We refer the reader to Appendix D for more details and examples.
@@ 3.2 Evaluation Metrics
We employ several forms of evaluation for ELM. First, given the inherent subjectivity of some tasks (e.g., movie reviews), human raters play a key role in gauging output quality. We ask 100 raters to assess ELM’s output w.r.t. its consistency with a movie’s plot, its linguistic coherence, and overall task quality. Each output is rated on a scale from 0 (completely irrelevant/incoherent) to 1 (highly relevant/coherent). The average rater’s score provides a holistic assessment of ELM’s performance.
@@
where projects the output to the domain embedding space and is the distance metric in (or alternatively, a similarity score); see Appendix A, Fig. 6 for an illustration.777In our experiments we use the same DLM to serve as , i.e., to re-embed output text into semantic space. The utility of SC is grounded in the expectation that if the generated output holds true to its source , then its embedded representation should not deviate substantially from its generating embedding vector. Hence, high SC suggests ELM has well-aligned domain embedding vectors with their (task-specific) descriptions. Importantly, this approach is not tied to examples in the data, providing the flexibility to measure the semantic consistency of hypothetical entities.
| Task (Appendix D) | Consistency Metrics | Human Evaluation (100 raters) | |||
@@
where are the ranks of the movies in for user provided by the language-based ranker, and are the (MovieLens) ground-truth ratings for provided by (converted to rankings); see Appendix A, Fig. 7 for an illustration.888In our experiments we use the dual-encoder DLM to serve as , by embedding test movies and user profiles and using their similarity for scoring/ranking. Here, RankingScore is any rank-comparison metric which compares to the ground-truth ranking . Examples include NDCG (Burges et al., 2005) and the Spearman rank-correlation coefficient. Finally, we measure behavioral consistency of movies in a fully analogous fashion (see Appendix A for a formal definition).
@@ 4 Empirical Analysis
We train two models: one (joint) model for the 24 movie tasks, and another for user profiles. For the movie tasks, we use 1000 randomly sampled examples for test and the rest for training. For the user profile task we use an 80/20 training/test split. Both models are fine-tuned using a pretrained PaLM 2-XS (Otter). We use a two-layer MLP for the domain-embedding adapter layer , and two-stage training, as described in Sec. 2.2. Particularly, for movie tasks, we run the first stage of training for 20k iterations, and then the second stage for another 300k iterations, using a batch size of 32 (i.e., roughly seven epochs over the whole dataset). We use a similar procedure for user profiles. We found training in two stages to be essential for convergence of ELM (see Appendix C for further discussion on two-stage training).
Test-item Evaluation. Table 1 shows consistency (SC and BC) and human rater results for all movie tasks on the test set. For each task, we ask 100 human raters to rate the consistency and comprehensiveness of our results w.r.t. the ground-truth data. We also ask qualitative task-specific questions (Appendix B describes rater instructions in detail). Rater evaluations show that ELM is able to generalize well across most tasks. Moreover ELM maintains high semantic consistency across all tasks and good behavioral consistency for most. We note that performance tends to be worse on two-movie tasks, due to the additional challenge of “conversing” about similarities and differences between two distinct embedding representations. We emphasize the importance of our consistency metrics, as they reflect how well ELM adapts to the domain embedding space .
@@
We first assess consistency metrics on interpolations of two movies or users. Fig. 3 compares the performance of ELM with state-of-the-art text-only LLMs; namely, PaLM 2-L and GPT 4. For movies, we linearly interpolate (with mixture coefficient ) embedding vectors of movies in the training set with the classic movie “Forrest Gump (1994)", and provide the interpolated embedding to ELM. For users, we interpolate random pairs of user embedding vectors from the training set. For a fairer comparison, we provide the text-only LLMs with a ground-truth examples of each task’s output, and prompt them to generate the task output. Particularly for users, we provide the text-only LLMs with ground-truth user profiles, and instruct them to interpolate between them (see Appendix E for the precise prompts used). We measure semantic consistency for movie tasks and behavioral consistency for the user profile task. We found that text-only LLMs are significantly less consistent than ELM. This is not surprising, as describing a domain embedding space manifold to a text-only LLM is especially challenging. Moreover, ELM is able to generate high-quality, consistent responses to our tasks. We provide further qualitative results in Appendix F.
We next test ELM’s consistency as we vary the degree of interpolation between movies, by varying the mixture weight . Fig. 4 shows semantic consistency (magenta line) for two of our movie tasks. We find that outputs are highly consistent across all degrees of interpolation values, but with a tendency for slightly lower consistency scores for embeddings that are farther away from the two movies. This result is unsurprising, as we expect our model to perform best on “real” data points, and degrade as it moves from these. Also note these are linear interpolations, whereas the actual embedding manifold may be highly non-linear.
@@
In Fig. 4, we see that semantic consistency of movie extrapolations generally decreased as we extrapolated further in CAV directions. This is expected as extrapolations outside the data manifold should induce higher error. Similarly, Fig. 5 shows behavioral consistency results for CAV extrapolations of user profiles. These results show attribute extrapolations to be behaviorally consistent for different values. Perhaps surprisingly, user profiles show greater behavior consistency as we increase the degree to which they move in a CAV direction. We hypothesize this is due to user profiles becoming “more predictable" in our CF model by making user preferences more extreme. For instance, as we make a user profile gravitate to funnier movies, preferences tend to degenerate with a near-exclusive focus on funny movies (thus rendering rating prediction – and especially relative rating prediction – more accurate). This in turn increases BC. We provide further qualitative results for interpolations and CAV-based extrapolations in Fig. 1. We refer to Appendix F for more qualitative results.
@@ 6 Conclusion
Future directions include: investigation of alternative architectures and training methods for interpreting domain embeddings (e.g., RLHF/RLAIF, see Appendix G), additional metrics for evaluation, and coupling ELM with analytic and optimization techniques for generating useful hypothetical embedding vectors.
Appendix A Semantic and Behavioral Consistency

### v2

@@ 1 Introduction
The success of deep learning has brought forth a paradigm shift in knowledge representation through the concept of embeddings—dense vector representations that capture high-dimensional information about entities, concepts, or relationships in a compact and useful format. Embeddings are ubiquitous, finding application in natural language processing (Mikolov et al., 2013; Devlin et al., 2018), recommender systems (Rendle, 2010), protein sequence modeling (Rives et al., 2021), and more. These embeddings are invaluable, capturing nuanced relationships and semantic structure in data which traditional machine learning (ML) approaches often miss. Nevertheless, understanding these abstract representations remains challenging.
By design, the structure and underlying information carried by an embedding is heavily adapted to the idiosyncrasies of the downstream task, posing a substantial challenge to its interpretation and manipulation. Previous work on ML interpretability offers various task-independent means to interpret embeddings, including dimensionality reduction techniques (e.g., t-SNE (Van der Maaten & Hinton, 2008), UMAP (McInnes et al., 2018)) or concept activation vectors (CAVs) (Kim et al., 2018). While very useful, these techniques are fairly narrow in scope.
As an alternative to such interpretability methods, suppose one could engage with embeddings using natural language to query information not directly expressed by the name or description of the underlying entity/concept. Perhaps one could even extract information from the embedding representations of non-existent or hypothetical entities. Consider, for instance, an embedding space representing items from an online commerce site, trained using a large corpus of user ratings or purchases, reviews, and other multi-faceted data sources. The embedding representation of an item may implicitly embody intricate details about its quality, usability, design, customer satisfaction, etc. Moreover, suppose one wanted to understand the properties of a hypothetical item at a specific point in the embedding space, say, if we predicted a potential market for such an item. Since no such item (or description) exists, asking a conventional language model to describe this embedding point would be futile. However, a large language model (LLM) trained to interpret the embedding representation itself could handle such a task (see Fig. 1 for an illustration).
@@ 2 Integrating Domain Embeddings with Large Language Models
We assume a domain embedding which maps entities in some vocabulary (e.g., a set of known users and items) into a latent space , equipped with a distance metric . The embedding vector of any is typically a representation used in some downstream task, and is trained to encompass the latent features of necessary for this task. For example, in a recommender system, collaborative filtering (CF) (Salakhutdinov & Mnih, 2007; Koren et al., 2011) is often used to generate embedding vectors of users and items s.t. the dot product (or cosine similarity) of and is predictive of user ’s affinity for item . In image classification, might generate embedding representations useful for object detection (Zou et al., 2023), while in search/information retrieval, might embed both queries and documents in to measure a document’s relevance to a query (Zuccon et al., 2015; Luan et al., 2021).
We assume a pretrained LLM, , which maps (length ) sequences of language tokens into another. Modern LLMs are composed of two parts, . The first part is an embedding layer, (not to be confused with the domain embeddings ), which maps tokens to their respective token embedding representations (as distinct from the domain embedding). The second part is the the dense model, , which maps a sequence of token embeddings to a sequence of tokens 222In practice, the dense model outputs logits over tokens, which are sampled by a decoding procedure. (e.g., transformer, Vaswani et al. (2017)).
@@
Domain embedding space | Domain space metric | ||
Token (language) embedding space | Token space | ||
LLM embedding | Pretrained LLM | ||
Adapter () | Embedding Language Model | ||
Domain embedding | LLM excluding input embedding layer |
@@
Training Procedure. Given a downstream loss function , we can differentially optimize the model with parameters over the tasks in by solving . Aside from maximum likelihood-based training, we discuss optimizing using reinforcement learning from AI feedback (Lee et al., 2023) in Appendix H.
Training continuous prompts poses challenges due to constraints over the pretrained token embeddings, which may result in convergence to local minima (Liu et al., 2021). Intuitively, as the pretrained embedding layer readily maps to language space, the newly initialized embedding layer may require numerous updates to map to a similar space. To overcome this, we divide training into two stages. In the first stage, we train the adapter on tasks in by keeping all other parameters (, ) frozen. Since is pretrained, the learned first-stage mapping from to improves convergence in the next stage. In the second stage, we fine-tune the full model by training all parameters (, , ). To increase efficiency, we can update a smaller number of parameters in alternating fashion (Hu et al., 2021). We found that two-stage training is essential for convergence of (see Appendix D for further discussion of the two-stage training procedure).
@@ 3.1 Data
Training Data and Tasks. We test two versions of ELM, one that interprets semantic embeddings of movies, and a second which interprets behavioral embeddings of users. For the first model, we use a variety of tasks to test the model’s ability to extract reasonable semantic information from the domain embeddings. More specifically, we construct 24 movie-focused tasks using a pretrained PaLM 2-L model. For each task, we generate training data by prompting the PaLM 2-L with a movie’s title and additional task-specific information (e.g., writing review, listing characteristics, comparing movies, see details below). We then use PaLM 2-L’s generated output as training targets for ELM. As for training inputs, we provide ELM with the same prompts used in PaLM 2-L, where the title of the movie is replaced by its semantic embedding . We emphasize that ELM does not receive any information about the movie (including its title), apart from its semantic embedding vector.
Our 24 movie tasks include single movie semantic tasks, such as describing a movie plot or summarizing a movie; single movie subjective tasks, such as writing positive or negative reviews for a movie; and movie pair subjective tasks, such as comparing characteristics of movies. Appendix E provides a complete description of all 24 tasks, including the prompts used, and sample outputs.
Finally, we train an additional model to interpret user behavioral embeddings, by generating user preference profiles. For this task, we generate a textual summary of a user’s preferences as follows: (a) we sample five positively rated and five negatively rated movies for each user in MovieLens 25M; (b) we then prompt PaLM 2-L to describe (in ten bullet points) characteristics of a person who likes the first five movies, but dislikes the second five. This resulting summary is used as training output for ELM on the user-profile generation task. The resulting task is to generate such a user profile from a user’s behavioral embedding (i.e., interpret the domain embedding). Similar to the movie tasks, no other information about the user, apart from its embedding, is provided to ELM. We refer the reader to Appendix E for more details and examples.
@@ 3.2 Evaluation Metrics
We employ several forms of evaluation for ELM. First, given the inherent subjectivity of some tasks (e.g., movie reviews), human raters play a key role in gauging output quality. We ask raters to assess ELM’s output w.r.t. its consistency with a movie’s plot, its linguistic coherence, and overall task quality. Each output is rated on a scale from 0 (completely irrelevant/incoherent) to 1 (highly relevant/coherent). The average rater’s score provides a holistic assessment of ELM’s performance.
@@
where projects the output to the domain embedding space and is the distance metric in (or alternatively, a similarity score); see Appendix B, Fig. 6 for an illustration.666In our experiments we use the same DLM to serve as , i.e., to re-embed output text into semantic space. The utility of SC is grounded in the expectation that if the generated output holds true to its source , then its embedded representation should not deviate substantially from its generating embedding vector. Hence, high SC suggests ELM has well-aligned domain embedding vectors with their (task-specific) descriptions. Importantly, this approach is not tied to examples in the data, providing the flexibility to measure the semantic consistency of hypothetical entities.
Task (Appendix E) | Consistency Metrics | Human Evaluation (100 raters) | |||
@@
where are the ranks of the movies in for user provided by the language-based ranker, and are the (MovieLens) ground-truth ratings for provided by (converted to rankings); see Appendix B, Fig. 7 for an illustration.777In our experiments we use the dual-encoder DLM to serve as , by embedding test movies and user profiles and using their similarity for scoring/ranking. Here, RankingScore is any rank-comparison metric which compares to the ground-truth ranking . Examples include NDCG (Burges et al., 2005) and the Spearman rank-correlation coefficient. Finally, we measure behavioral consistency of movies in a fully analogous fashion (see Appendix B for a formal definition).
@@ 4 Empirical Analysis
We train two models: one (joint) model for the 24 movie tasks, and another for user profiles. For the movie tasks, we use 1000 randomly sampled examples for test and the rest for training. For the user profile task we use an 80/20 training/test split. Both models are fine-tuned using a pretrained PaLM 2-XS (Otter). We use a two-layer MLP for the domain-embedding adapter layer , and two-stage training, as described in Sec. 2. Particularly, for movie tasks, we run the first stage of training for 20k iterations, and then the second stage for another 300k iterations, using a batch size of 32 (i.e., roughly seven epochs). We use a similar procedure for user profiles. We found training in two stages to be essential for convergence of ELM (see Appendix D for further details).
Test-item Evaluation. Table 2 shows consistency (SC and BC) and human rater results for all movie tasks on the test set. For each task, we ask 100 human raters888Raters were paid contractors. They received their standard contracted wage, which is above the living wage in their country of employment. to rate the consistency and comprehensiveness of our results w.r.t. the ground-truth data. We also ask qualitative task-specific questions (Appendix C describes rater instructions in detail). Rater evaluations show that ELM is able to generalize well across most tasks. Moreover ELM maintains high semantic consistency across all tasks and good behavioral consistency for most. We note that performance tends to be worse on two-movie tasks, due to the additional challenge of “conversing” about similarities and differences between two distinct embedding representations. We emphasize the importance of our consistency metrics, as they reflect how well ELM adapts to the domain embedding space .
@@
We first assess consistency metrics on interpolations of two movies or users. Fig. 3 compares the performance of ELM with state-of-the-art text-only LLMs; namely, PaLM 2-L and GPT 4. For movies, we linearly interpolate (with mixture coefficient ) embedding vectors of movies in the training set with the classic movie “Forrest Gump (1994)", and provide the interpolated embedding to ELM. For users, we interpolate random pairs of user embedding vectors from the training set. For a fairer comparison, we provide the text-only LLMs with a ground-truth examples of each task’s output, and prompt them to generate the task output. Particularly for users, we provide the text-only LLMs with ground-truth user profiles, and instruct them to interpolate between them (see Appendix F for the precise prompts used). We measure semantic consistency for movie tasks and behavioral consistency for the user profile task. We found that text-only LLMs are significantly less consistent than ELM. This is not surprising, as describing a domain embedding space manifold to a text-only LLM is especially challenging. Moreover, ELM is able to generate high-quality, consistent responses to our tasks. We provide further qualitative results in Appendix G.
We next test ELM’s consistency as we vary the degree of interpolation between movies, by varying the mixture weight . Fig. 4 shows semantic consistency (magenta line) for two of our movie tasks. We find that outputs are highly consistent across all degrees of interpolation values, with a tendency for lower consistency scores for embeddings that are farther away from the two movies. This result is unsurprising, as we expect performance to degrade as it moves away from “real" data points. Also note these are linear interpolations, though the true embedding manifold may be highly non-linear.
@@
In Fig. 4, we see that semantic consistency of movie extrapolations generally decreased as we extrapolated further in CAV directions. This is expected as extrapolations outside the data manifold should induce higher error. Similarly, Fig. 5 shows behavioral consistency results for CAV extrapolations of user profiles. These results show attribute extrapolations to be behaviorally consistent for different values. Perhaps surprisingly, user profiles show greater behavior consistency as we increase the degree to which they move in a CAV direction. We hypothesize this is due to user profiles becoming “more predictable" in our CF model by making user preferences more extreme. For instance, as we make a user profile gravitate to funnier movies, preferences tend to degenerate with a near-exclusive focus on funny movies (thus rendering rating prediction – and especially relative rating prediction – more accurate). This in turn increases BC. We provide further qualitative results for interpolations and CAV-based extrapolations in Fig. 1. We refer to Appendix G for more qualitative results.
@@ 6 Conclusion
Appendix A Additional Experiments on the Amazon Product Dataset
We complement our experiments and demonstrate ELM on the public Amazon dataset (mcauley2015image), which consists of 9.35M items with textual descriptions, 20.9M users, 233.1M reviews, and 82.83M ratings. We focus specifically on a subset of this dataset in the category “Clothing, Shoes and Jewelry", which consists of 5.7M reviews for 1.5M products.
Domain Embeddings.
Similar to our setting in the Movielens dataset, we create both semantic and behavioral embeddings. We train behavioral embeddings on users’ rating data using the same method described in Sec. 3.1. For semantic embeddings, we use DLM (see Sec. 3.1) to encode a concatenation of the item title, item description, item categories, and item features (which include item tags).
Training Data and Tasks.
We test ELM’s ability to interpret semantic and behavioral embeddings on a set of six tasks: item description, positive review, negative review, neutral review, item endorsements, and user profile. Item descriptions are taken directly from the Amazon dataset. For review data, we split the Amazon reviews w.r.t. to positive, negative, and neutral reviews. To ensure the datasets are of comparable size, we use scores of 1-3 to represent negative reviews, a score of 4 to represent neutral reviews, and a score of 5 to represent positive reviews. To create the item endorsement task we prompt a PaLM 2-L with the item title and description and ask it to create an endorsement for that product. Finally, for user profiles, we use a similar procedure as in Movielens. Specifically, we describe a user using positive and negative reviews they have posted, and ask a PaLM 2-L to describe a user with such reviews in ten bullet points. We note that, similar to the Movielens tasks, the Amazon training data does not consist of any information about items or users except for the behavioral or semantic embedding information. Table 3 summarizes the set of tasks we use.
Task (Appendix E) | Consistency Metrics | Human Evaluation (100 raters) | |||
Semantic (Cosine) | Behavioral (Rank Corr.) | Consistency with Item / User | Language Comprehensiveness | Task Quality | |
description | 0.93 | 0.92 | 0.81 | 0.88 | 0.75 |
positive review | 0.95 | 0.89 | 0.76 | 0.79 | 0.93 |
negative review | 0.91 | 0.94 | 0.77 | 0.76 | 0.71 |
neutral review | 0.96 | 0.82 | 0.8 | 0.79 | 0.7 |
endorsement | 0.9 | 0.81 | 0.91 | 0.93 | 0.94 |
user profile | 0.89 | 0.87 | 0.84 | 0.77 | 0.75 |
Results.
Table 4 shows consistency (SC and BC) and human rater results for the Amazon dataset tasks on a test set of 1000 examples (randomly selected). For each task, we ask 100 human raters to rate various aspects of the outputted result, such as its consistency with the true item, comprehensiveness of the language, and other task-specific quality evaluations. Simliar to the Movielens experiments, our results show that ELM is able to generalize well across all tasks. Moreover ELM maintains high semantic and behavioral consistency across all tasks.
We conduct further detailed human evaluation studies on the Amazon tasks. For this evaluation we ask human raters to evaluate both ground-truth data as well as results generated by our model. For example, we ask human raters to evaluate the quality of a review in the real data w.r.t. the ground-truth item description. This evaluation enables us to obtain a baseline for the model’s scores. Below we present examples of forms that were presented to the raters, with task-specific questions. In Tables 6 and 5 we show the average human rater answers to these questions. We find that our generated results correctly correspond to test items, and perhaps surprisingly, often surpasses the original ground-truth data quality, as viewed by the human raters.
Task | Q1 | Q2 | Q3 | Q4 | Q5 | Q6 | Q7 | Q8 | Q9 | Q10 |
description | 0.77 | 0.7 | 0.86 | 0.71 | 0.73 | 0.7 | 0.81 | 0.9 | 0.9 | 0.9 |
Task | Q1 | Q2 | Q3 | Q4 |
positive review | 0.91 | 0.92 | 0.92 | 0.98 |
negative review | 0.75 | 0.72 | 0.71 | 0.75 |
neutral review | 0.85 | 0.8 | 0.79 | 0.7 |
Task | Q1 | Q2 | Q3 |
endorsement | 0.91 | 0.93 | 0.93 |
Task | Q1 | Q2 | Q3 | Q4 | Q5 | Q6 | Q7 | Q8 | Q9 | Q10 |
description | 0.68 | 0.64 | 0.85 | 0.83 | 0.82 | 0.83 | 0.84 | 0.74 | 0.73 | 0.73 |
Task | Q1 | Q2 | Q3 | Q4 |
positive review | 0.85 | 0.76 | 0.79 | 0.92 |
negative review | 0.75 | 0.77 | 0.74 | 0.71 |
neutral review | 0.84 | 0.79 | 0.81 | 0.7 |
Below we provide several examples of user profiles in the Amazon dataset, as generated from positive and negative reviews. We used these profiles as training data for ELM to generalize to unseen user embeddings.
Appendix B Semantic and Behavioral Consistency